OpenStudy (chaylaceyx3):

A reaction will only be spontaneous at low temperatures if both delta H and delta S are negative. For a reaction in which delta H= -330.1 kJ/mol and delta S= -98.00 J/Kmol, determine the temperature (in degrees C) below which the reaction is spontaneous.

OpenStudy (aaronq):

a spontaneous reaction must have \(\Delta G>0\) thus \(\Delta G=\Delta H - T \Delta S\rightarrow 0=\Delta H - T\Delta S\) Solving for temp gives: \(T=\dfrac{\Delta H}{\Delta S}\)

OpenStudy (chaylaceyx3):

I got 3.37 and its saying the answer ins wrong @aaronq

OpenStudy (aaronq):

\(T=\dfrac{ -330.1~ kJ/mol }{98.00 ~J/Kmol}\) convert units of energy, 1 kJ = 1000 J \(T=\dfrac{ -330100~ J/mol }{98.00 ~J/Kmol}=3368.3673~K\)